OpenStudy (spiffy):

What is the equation of the following line? Be sure to scroll down first to see all answer options.

OpenStudy (1davey29):

I don't see any answer choices, but to solve this, first find your slope of the line.

OpenStudy (spiffy):

A. y=-6x B. y=6x C. y=-1/2x D. y=1/2x E. y=3x F. y=2x

OpenStudy (mathmale):

The illustration shows 2 points on the line. I'd suggest you begin by finding the slope of the line thru these 2 points. This is an essential skill. Then, using this slope and one of the points, write the equation of this line in point-slope form. Alternatively: Find the y-intercept of this line by inspection. Insert your slope and this y-intercept into the equation y=mx+b.

likeabossssssss (likeabossssssss):

hello welcome to os :)

OpenStudy (mathmale):

It'd be Spiffy if you'd please show your calculation of the slope, m.